How old am I if 400 reduced by 2 times my age is 244?
Answer by Alan3354(69443)

How old am I if 400 reduced by 2 times my age is 244?
-----------
400 - 2*Age = 244
-2*Age = -156
Age = 78 years, months, days, hours. No units were specified.